मराठी
तामिळनाडू बोर्ड ऑफ सेकेंडरी एज्युकेशनएचएससी वाणिज्य इयत्ता १२

Prove that V(X + b) = V(X)

Advertisements
Advertisements

प्रश्न

Prove that V(X + b) = V(X)

बेरीज
Advertisements

उत्तर

L.H.S = V(x + b)

= E(x + b)2 – [E(x + b)]2

= E(x2 + 2bx + b2) – [E(x) + b]2 

= E(x2) + 2bE(x) + b2 – [E(x)]2 + b2 + 2bE(x)]

= E(x2) + 2bE(x) + b2 – [E(x)]2 + b2 – 2bE(x)]

= E(x2) – [E(x)]2

= V(x)

= R.H.S

L.H.S = R.H.S

Hence proved.

shaalaa.com
Mathematical Expectation
  या प्रश्नात किंवा उत्तरात काही त्रुटी आहे का?
पाठ 6: Random Variable and Mathematical expectation - Miscellaneous problems [पृष्ठ १४४]

APPEARS IN

सामाचीर कलवी Business Mathematics and Statistics [English] Class 12 TN Board
पाठ 6 Random Variable and Mathematical expectation
Miscellaneous problems | Q 8. (ii) | पृष्ठ १४४

संबंधित प्रश्‍न

For the random variable X with the given probability mass function as below, find the mean and variance.

`f(x) = {{:(1/10, x = 2","  5),(1/5, x = 0","  1","  3","  4):}`


A commuter train arrives punctually at a station every half hour. Each morning, a student leaves his house to the train station. Let X denote the amount of time, in minutes, that the student waits for the train from the time he reaches the train station. It is known that the pdf of X is

`f(x) = {{:(1/30, 0 < x < 30),(0, "elsewhere"):}`
Obtain and interpret the expected value of the random variable X


The time to failure in thousands of hours of an electronic equipment used in a manufactured computer has the density function
`f(x) = {{:(3"e"^(-3x), x > 0),(0, "eleswhere"):}`
Find the expected life of this electronic equipment


Choose the correct alternative:

If P(X = 0) = 1 – P(X = 1). If E[X] = 3 Var(X), then P(X = 0) is


Define Mathematical expectation in terms of discrete random variable


A person tosses a coin and is to receive ₹ 4 for a head and is to pay ₹ 2 for a tail. Find the expectation and variance of his gains


Choose the correct alternative:

Demand of products per day for three days are 21, 19, 22 units and their respective probabilities are 0.29, 0.40, 0.35. Profit per unit is 0.50 paisa then expected profits for three days are


Choose the correct alternative:

A probability density function may be represented by


Choose the correct alternative: 

`int_(-oo)^oo` f(x) dx is always equal to


What is the expected value of a game that works as follows: I flip a coin and if tails pay you ₹ 2; if heads pay you ₹ 1. In either case, I also pay you ₹ 0.50


Share
Notifications

Englishहिंदीमराठी


      Forgot password?
Use app×